There are 3 blue and 5 red balloons, what are the odds that you pick one blue balloon?
Answer by jim_thompson5910(35256)

P(blue) = n(blue)/n(total) = 3/(5+3)=3/8=0.375

note: n(blue) means "number of blue"

So there's a 37.5% chance that you'll select a blue balloon.